# Rigidity transition in polydisperse shear-thickening suspensions

### v1

@@ Abstract
Shear thickening suspensions of non-Brownian polydisperse particles are simulated in 2D using a discrete element method based algorithm (LF-DEM) at high packing fractions ($\phi$) and large non-dimensional stresses ($\sigma$). Rigidity analysis of the stress induced particle clusters is carried out using pebble game algorithm for polydisperse suspensions and compared with the statistically equivalent bidisperse systems. A critical value of the packing fraction, $\phi_{c}$, close to the shear jamming transition, $\phi_{J}^{\mu}$, ($\phi_{c}<\phi_{J}^{\mu}$) is obtained where rigid particle clusters begin to grow sharply. The growth is found to be characterized by a critical transition of an order parameter ($f_{\text{rig}}$), defined by the fraction of rigid particles which scales as, $f_{\text{rig}}\sim(\phi-\phi_{c})^{\beta}$ for $\phi>\phi_{c}$, and by the susceptibility scaling, $\chi_{\text{rig}}\sim|\phi-\phi_{c}|^{-\gamma}$, with exponents $\beta=1/8$ and $\gamma=7/4$, which are consistent with the critical exponents in 2D Ising model. The variations of $f_{\text{rig}}$ and $\chi_{\text{rig}}$ in polydisperse suspensions are found to be identical to that of the statistically equivalent bidisperse suspensions. Finite size scaling analysis shows a divergence of correlation length near $\phi_{c}$ following critical exponent $\nu=1.64$, consistent with the 2D continuum percolation transition. Furthermore, $\phi_{c}$ and $\phi_{J}^{\mu}$ are found to vary non-monotonically with polydispersity index and depend on the particle stiffness.
@@ I Introduction
The flow behaviour of dense suspensions of non-Brownian particles has attracted the attention of researchers for the past several decades because of the rich physics involved in these apparently simple systems and for their wide range of industrial applications such as, in chocolate refining, concrete mixing, ceramic processing, rocket fuel, etc [7, 40, 5, 17]. It is well known that in shear flow, depending on the particle volume fraction, applied stress, and particle interaction forces, dense suspensions show rate-dependent rheology such as, continuous shear thickening (CST) and discontinuous (DST) shear thickening [34, 61, 35, 38, 13]. Furthermore, depending on the strength of the externally applied shear stress ($\hat{\sigma}$) and frictional contact force between the particles (with $\mu$ being the static friction coefficient), dense suspensions undergo shear jamming at a solid volume fraction ($\phi_{J}^{\mu}$) much lower than the maximum random close packing fraction or frictionless jamming fraction ($\phi_{J}^{0}$[7, 34, 38, 46]. At sufficiently high solid fraction, close to shear jamming, the transition from CST to DST is explained by the increase in the number of stress-induced frictional contacts [54, 61, 27]. Typically, there exists a characteristic stress $\sigma_{0}\approx F_{0}/a^{2}$, proportional to the repulsive force scale $F_{0}$ acting between particles of characteristic radius $a$, which determines the relative strength of the externally applied stress and controls the transition from frictionless ($\phi_{J}^{0}$) to frictional jamming ($\phi_{J}^{\mu}$[61, 34, 55]. In the limit of low stress ($\hat{\sigma}\ll\sigma_{0}$), the contacts between the particles are mostly dominated by lubrication interaction leading to frictionless jamming at sufficiently high particle volume fraction. Whereas, in the high stress limit ($\hat{\sigma}\gg\sigma_{0}$), lubrication between the particles is overpowered by the frictional contact leading to frictional jamming. This transition in the flow property has been described by a cross-over scaling which correlates the increase in frictional contacts with increasing applied stress [33, 50, 61]. Moreover, the flow state of a shear thickening dense suspension depends on the applied stress, solid volume fraction and strength of frictional interactions.
@@
In addition to the applied stress and strength of frictional contacts, the size distribution of the particles immensely influence the shear thickening and jamming transition in these suspensions [16, 23, 32, 56, 10, 20, 29]. For instance, at a fixed solid volume fraction $\phi$, bidisperse suspensions of any given ratio of large particle volume ($\phi_{\ell}$) to total particle volume fraction ($\zeta=\phi_{\ell}/\phi$), are found to show lower viscosity as compared to a mono-disperse suspension at the same condition [10, 43, 32]. Furthermore, at a constant $\zeta$, the viscosity is observed to monotonically decrease with increasing bidispersity ratio ($\delta=a_{\ell}/a_{s}$), where $a_{\ell}$ and $a_{s}$ are the large and small particle size, respectively [16, 43]. This indicates more efficient packing of the particles in bidisperse suspensions as compared to monodisperse systems. Similar trends were observed experimentally for bidisperse suspensions of polymethylmethacrylate (PMMA) particles [23]. However, the Wyart and Cates (WC) model [61], used to explain the CST and DST in dense suspensions, failed to predict the shear thickening behaviour for PMMA suspensions with a predominant fraction of large particles [23]. The results suggest that the underlying physics in the shear thickening behaviour of dense suspensions also depends on the type of contacts between the particles, i.e. whether large-large, small-small or large-small contact is predominating. Recently, Malbranche et al. [32] simulated shear thickening suspensions with different values of bidispersity ratio ($\delta$) over a range of shear stress and observed the CST-DST transition to be dependent on the composition of small and large particles, supporting the observations by Guy et al. [23]. Nevertheless, they show large particle ordering in the low stress limit to explain a reduction in viscosity at high value of $\zeta(=0.85)$ and found that the Maron-Pierce (M-P) model [36] can fit the viscosity divergence with packing fraction both in the low and high stress limits [32]. Another important feature in bidisperse shear thickening suspensions is the non-monotonicity of the relative suspension viscosity ($\eta_{r}$) with $\zeta$ at a fixed bidispersity ratio $\delta$ [43, 32, 23, 10]. Several models have been proposed to explain this behaviour by correlating the maximum packing fraction $\phi_{J}$ with the particle size distribution [48, 49, 1, 44]. More recently, Singh et al. [56] proposed a model based on a micro-mechanical picture which indicates that the non-monotonicity in the relative viscosity is related to the non-monotonic behaviour of the stress contribution due to large-small particle contacts. Such a finding reiterates the important role of the types of particle contacts in affecting the rheology of bidisperse suspensions. While the bidisperse suspensions are the simplest form of polydisperse systems and are relatively easier to study, materials typically handled in industries and other practical problems generally consist of polydisperse particles with more than two particle sizes. However, due to the presence of multiple particle sizes, the effects of particle contact types on the rheology and flow microstructure are much difficult to analyse statistically. Interestingly, it has been shown that the rheology of a polydisperse suspension, in terms of viscosity, normal stresses, particle pressures and packing efficiency of the particles, is identical to that of a statistically equivalent bidisperse suspension [14, 41, 43, 22]. In case of dry granular materials, similar analogy between the flow behaviour of monodisperse and polydisperse systems is established by matching the grain size distribution [45]. However, while the bulk rheological properties of monodisperse and polydisperse shear thickening suspensions are well studied, a comprehensive understanding of the microstructural features leading to CST-DST transition and shear jamming in these systems is still missing. Specifically, the mechanisms involved in the growth of flow microstructure leading to a shear jammed state with increasing solid fraction, and how it is affected by particle size distribution is not well understood. Numerical analysis of the frictional contact networks and cluster rigidity suggests a correlation between huge stress fluctuations and percolation transition at the onset of DST and shear-jamming [53, 52, 21, 59]. Shear jamming in the high stress limit is found to be preceded by a non-equilibrium critical phenomenon described by a rigidity transition of frictionally contacted particle clusters [51]. While these studies are focused on understanding the correlation between the growth of the frictional contact network and DST or onset of shear jamming in nearly mono-disperse (or bidisperse suspensions with small $\delta$) suspensions, effects of polydispersity on the growth of flow microstructure is not well explored. In this work, we have studied the growth and transition of cluster rigidity near shear jamming in the high stress limit for systems of polydisperse suspensions.
@@
Recently, van der Naald et al. [59] and Santra et al. [51] have implemented a constraint counting algorithm (pebble game[25, 24, 28] to compute fraction of rigid clusters in two dimensional (2D) bidisperse suspensions to understand the underlying mechanism in shear jamming at high stresses. Interestingly, Santra et al. [51] reported a critical phenomenon associated with the shear jamming at large stresses based on an order parameter computed from the rigid particle fraction, with critical exponents similar to the 2D-Ising model. Inspired from these studies, here we have investigated the effects of particle size distribution on the formation of rigid clusters and the associated critical transition near shear jamming at large stress ($\hat{\sigma}\gg\sigma_{0}$). Specifically, we compare the viscosity, flow microstructure and gradient direction velocity correlation functions of polydisperse systems with the statistically equivalent bidisperse suspensions. Moreover, our analysis demonstrates that shear jamming in polydisperse suspensions is also associated with a critical phenomenon which is identical to the statistically equivalent bidisperse systems. We also perform finite size scaling to extract the critical exponents corresponding to the divergence of the correlation length near the critical transition and compare that with the 2D-Ising model and percolation transition. Broadly, we divided the rest of the paper into the following sections. Section II discusses the modeling and simulation methodology. The method of analysis and validation study for polydisperse systems are presented in section III. The key results are discussed in section IV and the main conclusions are highlighted in section V.
@@ II Modeling and simulations
In the present study we have simulated systems of polydisperse suspensions of non-Brownian particles in 2D, considering a monolayer of $N$ spherical particles in a 2D simulation box. It is well established that 2D and 3D suspensions essentially show similar bulk rheological behaviour with proper scaling of the particle packing fraction [18, 39, 51]. However, we have chosen a 2D system for this study to implement the pebble game algorithm for the analysis of rigid clusters, which is only applicable in 2D as discussed in the subsequent sections. The simulations are carried out using the well-established Lubrication Flow - Discrete Element Method (LF-DEM) algorithm developed by Mari et al. [34], Seto et al. [54]. The simulation methodology involved modeling of neutrally buoyant non-Brownian spherical particles with specified particle size distribution suspended in a Newtonian solvent, where the particles interact via pairwise lubrication interaction, frictional contact force and short range electrostatic repulsion. Simple shear flow is implemented by using Lees-Edward periodic boundary conditions. The equation of motion for each particle is given by the following force and torque balances derived from the Langevin equation in the overdamped limit, neglecting the Brownian forces.
$\mathbf{F}_{h}\left(\mathbf{x},\mathbf{u}\right)+\mathbf{F}_{c}\left(\mathbf{x}\right)+\mathbf{F}_{r}\left(\mathbf{x}\right)=\mathbf{0}$
$\mathbf{T}_{h}\left(\mathbf{x},\mathbf{u}\right)+\mathbf{T}_{c}\left(\mathbf{x}\right)=\mathbf{0}$
@@
$\mathbf{F}_{h}=-\mathbf{R}_{FU}\left(\mathbf{x}\right)\cdot\left[\mathbf{u}-\mathbf{u}^{\infty}\left(\mathbf{x}\right)\right]+\mathbf{R}_{FE}\left(\mathbf{x}\right):\mathbf{E}^{\infty}$ (1)
@@
$\mathbf{F}_{c,n}^{(i,j)}$ $=$ $k_{n}h_{ij}\bm{n}_{ij}+\gamma_{n}\mathbf{U}_{n}^{(i,j)}$
$\mathbf{F}_{c,\tan}^{(i,j)}$ $=$ $k_{t}\,\xi^{(i,j)}$
where, $k_{n}$ and $k_{t}$ are the normal and tangential spring constants, respectively, with $k_{t}=0.9\,k_{n}$, $\gamma_{n}$ is the damping coefficient for the normal component, $\mathbf{U}_{n}^{(ij)}$ is the normal component of the relative velocity between $i^{th}$ and $j^{th}$ particles, and $\xi^{(ij)}$ is the tangential stretch. The non-sliding frictional contacts between the particles are satisfied by the Coulomb friction law $|\mathbf{F}_{c,\tan}^{(i,j)}|\leq\mu|\mathbf{F}_{c,n}^{(i,j)}|$ with an interparticle friction coefficient $\mu$. The short-range repulsive component of the force is implemented by an electrostatic repulsion model given by,
[…]
where $\lambda=0.05\,a$ is considered as the Debye length. Further details on the simulation model are available in the work by Mari et al. [34].
@@
where, $\eta_{0}$ is the solvent viscosity, $\phi$ is the particle packing fraction, $\sigma_{c}$ and $\sigma_{r}$ are the shear stresses corresponding to the contact force and repulsive force, respectively. These stresses are computed as $\sigma_{r,c}=V^{-1}(\mathbf{x}\mathbf{F}_{r,c}-\mathbf{R}_{SU}\cdot\mathbf{R}^{-1}_{FU}\cdot\mathbf{F}_{r,c})_{xy}$ and the hydrodynamic stress is evaluated as $\dot{\gamma}\eta_{H}=V^{-1}[(\mathbf{R}_{SE}-\mathbf{R}_{SU}\cdot\mathbf{R}^{-1}_{FU}\cdot\mathbf{R}_{FE}):\mathbf{E}^{\infty}]_{xy}$, where $V$ is the volume of the simulation domain, and the resistance matrices $\mathbf{R}_{SU}$ and $\mathbf{R}_{SE}$ determine the lubrication stresses from the particle velocity and rate of deformation, respectively.

### v2

@@ Abstract
Shear thickening suspensions of non-Brownian polydisperse particles are simulated in 2D using a discrete element method based algorithm (LF-DEM) at high packing fractions ($\phi$) and large non-dimensional stresses ($\sigma$). Rigidity analysis of the stress induced particle clusters is carried out using pebble game algorithm for polydisperse suspensions and compared with the statistically equivalent bidisperse systems. A critical value of the packing fraction, $\phi_{c}$, close to the shear jamming transition, $\phi_{J}^{\mu}$, ($\phi_{c}<\phi_{J}^{\mu}$) is obtained where rigid particle clusters begin to grow sharply. The growth is found to be characterized by a critical transition of an order parameter ($f_{\text{rig}}$), defined by the fraction of particles in rigid clusters which scales as, $f_{\text{rig}}\sim(\phi-\phi_{c})^{\beta}$ for $\phi>\phi_{c}$, and by the susceptibility scaling, $\chi_{\text{rig}}\sim|\phi-\phi_{c}|^{-\gamma}$, with exponents having values consistent with the critical exponents in 2D percolation transition. The variations of $f_{\text{rig}}$ and $\chi_{\text{rig}}$ in polydisperse suspensions are found to be identical to that of the statistically equivalent bidisperse suspensions. Finite size scaling analysis shows a divergence of correlation length near $\phi_{c_{\infty}}$ following critical exponent $\nu\approx 1.33$, in agreement with the 2D percolation theory. Furthermore, $\phi_{c}$ and $\phi_{J}^{\mu}$ are found to vary non-monotonically with polydispersity index and depend on the particle stiffness.
@@ I Introduction
The flow behavior of dense suspensions of non-Brownian particles has attracted the attention of researchers for the past several decades because of the rich physics involved in these apparently simple systems and for their wide range of industrial applications such as, in chocolate refining, concrete mixing, ceramic processing, rocket fuel, etc [9, 45, 6, 18]. It is well known that in shear flow, depending on the particle volume fraction, applied stress, and particle interaction forces, dense suspensions show rate-dependent rheology such as, continuous shear thickening (CST) and discontinuous (DST) shear thickening [40, 68, 39, 43, 14]. Furthermore, depending on the strength of the externally applied shear stress ($\hat{\sigma}$) and frictional contact force between the particles (with $\mu$ being the static friction coefficient), dense suspensions undergo shear jamming at a solid volume fraction ($\phi_{J}^{\mu}$) much lower than the maximum random close packing fraction or frictionless jamming fraction ($\phi_{J}^{0}$[9, 40, 43, 51]. At sufficiently high solid fraction, close to shear jamming, the transition from CST to DST is explained by the increase in the number of stress-induced frictional contacts [60, 68, 32]. Typically, there exists a characteristic stress $\sigma_{0}\approx F_{0}/a^{2}$, proportional to the repulsive force scale $F_{0}$ acting between particles of characteristic radius $a$, which determines the relative strength of the externally applied stress and controls the transition from frictionless ($\phi_{J}^{0}$) to frictional jamming ($\phi_{J}^{\mu}$[68, 40, 62]. In the limit of low stress ($\hat{\sigma}\ll\sigma_{0}$), the contacts between the particles are mostly dominated by lubrication interaction leading to frictionless jamming at sufficiently high particle volume fraction. Whereas, in the high stress limit ($\hat{\sigma}\gg\sigma_{0}$), lubrication between the particles is overpowered by the frictional contact leading to frictional jamming. This transition in the flow property has been described by a cross-over scaling which correlates the increase in frictional contacts with increasing applied stress [37, 55, 68]. Moreover, the flow state of a shear thickening dense suspension depends on the applied stress, solid volume fraction and strength of frictional interactions.
@@
In addition to the applied stress and strength of frictional contacts, the size distribution of the particles immensely influence the shear thickening and jamming transition in these suspensions [17, 24, 38, 61, 11, 21, 34]. For instance, at a fixed solid volume fraction $\phi$, bidisperse suspensions of any given ratio of large particle volume ($\phi_{\ell}$) to total particle volume fraction ($\zeta=\phi_{\ell}/\phi$), are found to show lower viscosity as compared to a mono-disperse suspension at the same condition [11, 48, 38]. Furthermore, at a constant $\zeta$, the viscosity is observed to monotonically decrease with increasing bidispersity ratio ($\delta=a_{\ell}/a_{s}$), where $a_{\ell}$ and $a_{s}$ are the large and small particle size, respectively [17, 48]. This indicates more efficient packing of the particles in bidisperse suspensions as compared to monodisperse systems. Similar trends were observed experimentally for bidisperse suspensions of polymethylmethacrylate (PMMA) particles [24]. However, the Wyart and Cates (WC) model [68], used to explain the CST and DST in dense suspensions, failed to predict the shear thickening behavior for PMMA suspensions with a predominant fraction of large particles [24]. The results suggest that the underlying physics in the shear thickening behavior of dense suspensions also depends on the type of contacts between the particles, i.e. whether large-large, small-small or large-small contact is predominating. Recently, Malbranche,Nelya et al. [38] simulated shear thickening suspensions with different values of bidispersity ratio ($\delta$) over a range of shear stress and observed the CST-DST transition to be dependent on the composition of small and large particles, supporting the observations by Guy et al. [24]. Nevertheless, they show large particle ordering in the low stress limit to explain a reduction in viscosity at high value of $\zeta(=0.85)$ and found that the Maron-Pierce (M-P) model [41] can fit the viscosity divergence with packing fraction both in the low and high stress limits [38]. Another important feature in bidisperse shear thickening suspensions is the non-monotonicity of the relative suspension viscosity ($\eta_{r}$) with $\zeta$ at a fixed bidispersity ratio $\delta$ [48, 38, 24, 11]. Several models have been proposed to explain this behavior by correlating the maximum packing fraction $\phi_{J}$ with the particle size distribution [53, 54, 1, 49]. More recently, Singh et al. [61] proposed a model based on a micro-mechanical picture which indicates that the non-monotonicity in the relative viscosity is related to the non-monotonic behavior of the stress contribution due to large-small particle contacts. Such a finding reiterates the important role of the types of particle contacts in affecting the rheology of bidisperse suspensions. While the bidisperse suspensions are the simplest form of polydisperse systems and are relatively easier to study, materials typically handled in industries and other practical problems generally consist of polydisperse particles with more than two particle sizes. However, due to the presence of multiple particle sizes, the effects of particle contact types on the rheology and flow microstructure are much difficult to analyse statistically. Interestingly, it has been shown that the rheology of a polydisperse suspension, in terms of viscosity, normal stresses, particle pressures and packing efficiency of the particles, is identical to that of a statistically equivalent bidisperse suspension [15, 47, 48, 23]. In case of dry granular materials, similar analogy between the flow behavior of monodisperse and polydisperse systems is established by matching the grain size distribution [50]. However, while the bulk rheological properties of monodisperse and polydisperse shear thickening suspensions are well studied, a comprehensive understanding of the microstructural features leading to CST-DST transition and shear jamming in these systems is still missing. Specifically, the mechanisms involved in the growth of flow microstructure leading to a shear jammed state with increasing solid fraction, and how it is affected by particle size distribution is not well understood. Numerical analysis of the frictional contact networks and cluster rigidity suggests a correlation between huge stress fluctuations and percolation transition at the onset of DST and shear-jamming [58, 57, 22, 66]. Shear jamming in the high stress limit is found to be preceded by a non-equilibrium critical phenomenon described by a rigidity transition of frictionally contacted particle clusters [56]. While these studies are focused on understanding the correlation between the growth of the frictional contact network and DST or onset of shear jamming in nearly mono-disperse (or bidisperse suspensions with small $\delta$) suspensions, effects of polydispersity on the growth of flow microstructure is not well explored. In this work, we have studied the growth and transition of cluster rigidity near shear jamming in the high stress limit for systems of polydisperse suspensions.
@@
Recently, van der Naald et al. [66] and Santra et al. [56] have implemented a constraint counting algorithm (pebble game[28, 25, 33] to compute fraction of rigid clusters in two dimensional (2D) bidisperse suspensions to understand the underlying mechanism in shear jamming at high stresses. Notably, Santra et al. [56] reported a critical phenomenon associated with the shear jamming at large stresses based on an order parameter computed from the rigid particle fraction, with critical exponents similar to the 2D-Ising model. Inspired from these studies, here we have investigated the effects of particle size distribution on the formation of rigid clusters and the associated critical transition near shear jamming at large stress ($\hat{\sigma}\gg\sigma_{0}$). Specifically, we compare the viscosity, flow microstructure and gradient direction velocity correlation functions of polydisperse systems with the statistically equivalent bidisperse suspensions. Moreover, our analysis demonstrates that shear jamming in polydisperse suspensions is also associated with a critical phenomenon which is identical to the statistically equivalent bidisperse systems. We also perform finite size scaling to characterize the divergence of the correlation length and susceptibility near the critical transition and compare the critical exponents to that of 2D percolation transition. Broadly, we divided the rest of the paper into the following sections. Section II discusses the modeling and simulation methodology. The method of analysis and validation study for polydisperse systems are presented in section III. The key results are discussed in section IV and the main conclusions are highlighted in section V.
@@ II Modeling and simulations
In the present study we have simulated systems of highly concentrated polydisperse suspensions of non-Brownian particles in 2D, considering a monolayer of $N$ spherical particles confined on a plane in a simulation box. We have adapted this model for simulating 2D suspensions from the work by Seto et al. [59]. It is well established in the literature that the flow curves of shear thickening suspensions simulated in 2D and 3D are qualitatively similar [19, 44, 56]. In fact, the flow curves for 2D and 3D systems show nearly quantitative agreement if the packing fraction $\phi$ is appropriately mapped. However, we have chosen a 2D system for this study to implement the pebble game algorithm for the analysis of rigid clusters, which is only applicable in 2D as discussed in the subsequent sections. We agree that in 3D system the types of cluster would be topologically different from that observed in 2D, however, that does not result in any qualitative difference in the macroscopic flow properties. Thus 2D suspensions still show all the qualitative features of the flow behavior similar to 3D systems. The simulations are carried out using the well-established Lubrication Flow - Discrete Element Method (LF-DEM) algorithm developed by Mari,Romain et al. [40], Seto et al. [60]. The simulation methodology involved modeling of neutrally buoyant non-Brownian spherical particles with specified particle size distribution suspended in a Newtonian solvent, where the particles interact via pairwise lubrication interaction, frictional contact force and short range electrostatic repulsion. Simple shear flow is implemented by using Lees-Edward periodic boundary conditions. The LF-DEM simulation method is developed considering the flow to be in the regime of low Stokes number ($St\ll 1$) and particle Reynolds number, thereby neglecting the inertial term in the equations of motion. Furthermore, the size of the model particles are considered to be in the micron range where Brownian forces are unimportant. The long-range hydrodynamic interactions between the particles are assumed to be screened since the systems are highly concentrated and therefore, only pair-wise lubrication interactions are considered. In our simulations we have considered a constant microscopic friction coefficient where frictional states are only dependent of the applied stress. It is appropriate to mention here that while shear banding in concentrated dry granular media, which is a system very similar to the dense suspensions except for the hydrodynamic part, is an important physical phenomenon, in our study we do not observe it. This aspect is well studied in the literature and it has been found that for suspensions of non-Brownian particles with fixed microscopic friction, shear banding is generally not observed [31, 8, 26]. Moreover, LF-DEM provides an efficient and simplistic model for simulating systems of non-Brownian suspensions in the limit of high concentration. However, it is to be noted that the method will not work for dilute or semi-dilute suspensions where long-range hydrodynamic interactions become important. The method also requires significant modifications in case of simulating colloidal suspensions where Brownian forces are non-negligible. Considering all the assumptions discussed above, the equation of motion for each particle is given by the following force and torque balances derived from the Langevin equation in the overdamped limit, neglecting the Brownian forces.
$\mathbf{F}_{h}\left(\mathbf{x},\mathbf{u}\right)+\mathbf{F}_{c}\left(\mathbf{x}\right)+\mathbf{F}_{r}\left(\mathbf{x}\right)=\mathbf{0}$ (1)
$\mathbf{T}_{h}\left(\mathbf{x},\mathbf{u}\right)+\mathbf{T}_{c}\left(\mathbf{x}\right)=\mathbf{0}$ (2)
@@
$\mathbf{F}_{h}={}$ $-\mathbf{R}_{FU}(\mathbf{x})\cdot\bigl[\mathbf{u}-\mathbf{u}^{\infty}(\mathbf{x})\bigr]$
$+\mathbf{R}_{FE}(\mathbf{x}):\mathbf{E}^{\infty}$ (3)
@@
$\mathbf{F}_{c,n}^{(i,j)}$ $=$ $k_{n}h_{ij}\bm{n}_{ij}+\gamma_{n}\mathbf{U}_{n}^{(i,j)}$ (4)
$\mathbf{F}_{c,\tan}^{(i,j)}$ $=$ $k_{t}\,\xi^{(i,j)}$ (5)
where, $k_{n}$ and $k_{t}$ are the normal and tangential spring constants, respectively, with $k_{t}=0.9\,k_{n}$, $\gamma_{n}$ is the damping coefficient for the normal component, $\mathbf{U}_{n}^{(ij)}$ is the normal component of the relative velocity between $i^{th}$ and $j^{th}$ particles, and $\xi^{(ij)}$ is the tangential stretch. The non-sliding frictional contacts between the particles are satisfied by the Coulomb friction law $|\mathbf{F}_{c,\tan}^{(i,j)}|\leq\mu|\mathbf{F}_{c,n}^{(i,j)}|$ with an interparticle friction coefficient $\mu$. To mimic the stabilizing force observed in many experimental systems of non-Brownian suspensions, we have implemented a simplistic form of the electrostatic double-layer repulsive force given by,
$\mathbf{F}_{r}^{(ij)}(h)=\begin{cases}-2F_{0}\frac{a_{i}a_{j}}{a\,(a_{i}+a_{j})}\exp(-h/\lambda)\bm{n}_{ij}&\text{if $h_{ij}\geq 0$,}\\ -2F_{0}\frac{a_{i}a_{j}}{a\,(a_{i}+a_{j})}\bm{n}_{ij}&\text{if $h_{ij}<0$,}\end{cases}\quad$ (6)
where $\lambda=0.05\,a$ is considered as the Debye length. We would like to emphasize that we have used here an approximate form the double-layer electrostatic repulsion in the limit of small Debye length as compared to the particle radius [27, 40]. This approximation works well for the present study since the size of the non-Brownian particles are in the range of several microns and the corresponding Debye length $\lambda$ is of the order of $10^{-1}\,\mu m$. It should also be noted that the repulsive force is bounded to a maximum value since the particles are prevented from overlapping by the strong pairwise lubrication interaction with a regularization parameter $\delta$. Further details on the simulation model are available in the work by Mari,Romain et al. [40].
@@
where, $\eta_{0}$ is the solvent viscosity, $\phi$ is the particle packing fraction, $\sigma_{c}$ and $\sigma_{r}$ are the shear stresses corresponding to the contact force and repulsive force, respectively. These stresses are computed as,
$\sigma_{r,c}=V^{-1}(\mathbf{x}\mathbf{F}_{r,c}-\mathbf{R}_{SU}\cdot\mathbf{R}^{-1}_{FU}\cdot\mathbf{F}_{r,c})_{xy}$ (8)
The hydrodynamic stress is evaluated from the following relation
$\dot{\gamma}\eta_{H}=V^{-1}[(\mathbf{R}_{SE}-\mathbf{R}_{SU}\cdot\mathbf{R}^{-1}_{FU}\cdot\mathbf{R}_{FE}):\mathbf{E}^{\infty}]_{xy}$ (9)
where $V$ is the volume of the simulation domain, and the resistance matrices $\mathbf{R}_{SU}$ and $\mathbf{R}_{SE}$ determine the lubrication stresses from the particle velocity and rate of deformation, respectively.
Quantity dimensionless form characteristic scale
particle radius ($\hat{a}_{i}$) $a_{i}=\hat{a}_{i}/a$ $a=1.29\,\ell_{c}$
box size ($\hat{L}$) $L=\hat{L}/a$ $a=1.29\,\ell_{c}$
stress ($\hat{\sigma}$) $\sigma=\hat{\sigma}/\sigma_{0}$ $\sigma_{0}=F_{0}/6\pi a^{2}$
shear rate ($\dot{\gamma}$) $\dot{\gamma}/\dot{\gamma}_{0}$ $\dot{\gamma}_{0}=F_{0}/6\pi\eta_{0}a^{2}$
reduced viscosity ($\eta_{r}$) $\eta_{r}=\eta/\eta_{0}$ solvent viscosity ($\eta_{0}$)
time ($\hat{t}$) $t=\hat{t}\,\dot{\gamma}_{0}$ $1/\dot{\gamma}_{0}$
spring stiffness ($k_{n}$) $k_{n}/6\pi\eta_{0}a\dot{\gamma}$ $6\pi\eta_{0}a\dot{\gamma}$
[…]
All the simulations, if not mentioned otherwise, are carried out with $N=1000$ spherical particles, interacting with static friction coefficient $\mu=0.5$ at contact. The parameters and quantities used in the present study are all defined in terms of reduced units and a summary of this is provided in Table 1. It is to be noted that $\ell_{c}$, $F_{0}$ and $\eta_{0}$ are the fundamental quantities in our simulations and all other parameters, expressed in reduced units, are defined in terms of these fundamental quantities. In the present work, the quantity $\ell_{c}\,\gg 1\,\mu m$ is considered as the fundamental length unit and all the particle radii defined in terms of $\ell_{c}$ could be assumed to be well above $1\,\mu m$, resulting the suspensions to be non-Brownian. We have studied the systems in the limit of high non-dimensional stress ($\sigma=\hat{\sigma}/\sigma_{0}$) values of 25 and 100, where shear jamming is nearly independent of the applied stress [55, 62]. For consistency, all the length scales in the simulations are non-dimensionalized by a characteristic particle radius $a=1.29\,\ell_{c}$. To simulate polydisperse suspensions we have considered normal and log-normal particle size distributions. The mean ($\langle a\rangle$), standard deviation ($s.d.$) and skewness ($s$) of the distributions are defined as follows,
$\langle a\rangle$ $=$ $\displaystyle\frac{1}{N}\sum\limits_{i}a_{i}$ (10)
$s.d.$ $=$ $\left(\displaystyle\frac{1}{N}\sum\limits_{i}a_{i}^{2}-\langle a\rangle^{2}\right)^{1/2}$ (11)
$s$ $=$ $\left(\displaystyle\frac{1}{N}\sum\limits_{i}(a_{i}-\langle a\rangle)^{3}\right)/s.d.^{3}$ (12)
Typically, we refer to polydisperse system by its polydispersity index $\alpha$, defined as $\alpha=s.d./\langle a\rangle$. The details of the particle size distributions simulated in the present study is tabulated in Table 2.
@@
III Rigidity analysis and validation study
Refer to caption Refer to caption (a) (b)
Figure 1: (a) Schematic diagram illustrating a single step of the (3,3) PG algorithm. (b) Simulation snapshot showing regions of rigid clusters identified by PG algorithm (indicated by red coloured particles), surrounded by non-rigid regions (with blue coloured regions representing particles in frictional contacts and grey zones representing particles which are interacting via non-frictional interactions).
The identification of the rigid clusters from a given simulation snapshot is performed by applying the (3,3) pebble game (PG) algorithm, which is a constraint counting algorithm to identify the rigid nodes from a given graph [28, 42]. Here, each particle is assigned with 3 pebbles corresponding to 2 translational and 1 rotational degrees of freedom. In a contact network, whenever a pair of particles are in frictional contact, one of the particle loses one or two pebbles, depending on whether the contact is sliding or non-sliding, respectively, resulting in addition of constraints. We have illustrated this algorithm schematically in Fig. 1 (a). Likewise, for a given cluster of particles, all such constraints are constructed and the number of free pebbles are counted. If only a total of three free pebbles remain for the cluster, then the cluster is designated as a rigid one according to PG algorithm. It is to be noted that the rigid clusters are essentially transient structures identified by the PG algorithm, purely based on geometric criterion. The constraints implemented between the nodes (i.e. between the particles) are only dependent on the states of frictional contact, where frictionless hydrodynamic and repulsive interactions are not accounted for. Thus, by rigidity we imply the mechanical stability of a sub-structure of frictional contact network surrounded by non-rigid flowable regions, which are continuously reforming the clusters. This is illustrated schematically by presenting a simulation snapshot in Fig. 1 (b), where we have highlighted regions of rigid clusters (represented by red colored particles) surrounded by non-rigid regions (represented by blue and gray particles). Identification and growth of such rigid clusters have been considered in characterizing jamming in dry granular media [25, 16, 2] and also in understanding the mechanism of shear jamming in dense suspensions [66, 56]. In this work we have investigated the effects of polydispersity on the critical transition of cluster rigidity in terms of an order-parameter defined by the fraction of particles in rigid cluster, $f_{\text{rig}}$, same as that considered by Santra et al. [56] in there study on bidisperse systems.
$f_{\text{rig}}=\langle\frac{1}{N}\sum\limits_{i=1}^{N}n_{i}\rangle$ (13)
$\chi_{\text{rig}}=N\langle\frac{1}{N}\sum\limits_{i=1}^{N}(n_{i}-f_{\text{rig}})^{2}\rangle$ (14)
In the subsequent section we discuss the scaling behavior of $f_{\text{rig}}$ and $\chi_{\text{rig}}$ for the polydisperse systems and compared that with the bidisperse cases.